Please help find the inverse of f(x)= -3/(x+1)
Thanks Alot!!

Answer by nerdybill(7384)

Rewriting:
f(x)= -3/(x+1)
as
y = -3/(x+1)
.
Now, swap x and y:
x = -3/(y+1)
Solve for y:
x(y+1) = -3
y+1 = -3/x
y = -3/x -1
.
So, the inverse is:
f^(-1)(x) = -3/x -1